### v4

# Accurate Small-Signal Modeling of Digitally Controlled Buck Converters with ADC-PWM Synchronization

Abstract
Digital control has become increasingly widespread in modern power electronic converters. When acquiring feedback signals such as the inductor current, synchronizing the analog-to-digital converter (ADC) with the digital pulse-width modulator (DPWM) is commonly employed to accurately track their steady-state average. However, the small-signal implications of such synchronization have not been investigated. This paper presents an exact small-signal model for digitally controlled buck converters operating in forced continuous-conduction mode (FCCM) under constant-frequency current-mode control, explicitly accounting for ADC-DPWM synchronization. Using a sampled-data framework, the proposed model captures all sideband effects introduced by the sampling process, yielding precise predictions of both analog and digital loop gains, even at frequencies beyond the switching and sampling frequencies. Both asymmetrical and symmetrical carrier modulations are considered. Furthermore, the digital loop gain is derived in closed form using the modified z-transform, enabling low-complexity compensator design and stability assessment. Within this framework, the analog loop gain can be directly obtained from the digital loop gain, thereby eliminating the need for computationally intensive infinite series evaluations. The validity of the proposed model is confirmed through both simulation and experimental results.
I Introduction
As digital control becomes increasingly popular in power converters, its advantages in cost, programmability, and integration compare to analog designs have made it the preferred choice. [1] In particular, it greatly simplifies the realization of sophisticated algorithms. Moreover, it enables over-the-air (OTA) firmware updates, allowing deployed products to gain new functionality without hardware modifications.
In common engineering practice, digital power converter design typically begins with power stage modeling using the classical state-space averaging (SSA) technique [2, 3]. An analog compensator is then designed in the continuous-time domain, discretized, and finally implemented in the digital controller. By averaging the piecewise linear state-spaces over one switching period, SSA eliminates the non-linearities. However, SSA neglects essential effects, such as computational delays, sampling, the presence of ripple, and sideband component coupling [4]. Consequently, SSA loses validity even if the perturbation frequency is far below the switching frequency. While SSA leverages well-established linear time-invariant (LTI) control theory for rapid design, it fails to provide an accurate prediction of the closed-loop stability, particularly when the crossover frequency approaches the sampling frequency, leading to a potential risk of instability.
To extend the model accuracy near or beyond the switching frequency, the sideband effect must be taken into account to form a multifrequency model[5]. As an extension to SSA, the generalized state-space averaging method (GSSA)[6, 7, 8, 9] introduces Fourier series to account for higher-order harmonics in the time-domain waveforms. Based on the linear time-periodic (LTP) theory[10], the harmonic state-space (HSS)[11, 12] and harmonic transfer function (HTF)[13, 14] methods have been proposed to model frequency coupling effects. HSS employs the harmonic balance principle[15] to represent an LTP system as an infinite-dimensional matrix equation set[5], while HTF captures both direct and cross-frequency interactions. While these multifrequency modeling methods greatly enhance accuracy, they tend to produce high-order models that are computationally intensive. Consequently, computer-based tools are required, and obtaining a simple closed-form solution is often not possible.
As a compromise between model accuracy and computational complexity, researchers have proposed simplified models that include only a limited number of sidebands. Two-frequency models proposed by [16] and [17] considered one sideband component other than the perturbed frequency $f_{p}$. By considering a pair of sideband components and the switching frequency component, [18] fixes the low-frequency phase error in two-frequency models. An extended frequency model that incorporates all sideband components to enhance accuracy under large ripple conditions was proposed in [19]; However, the analytical form of the transfer function was not provided.
The describing function (DF) method [20, 21] enables accurate derivation of closed-loop transfer functions and has been successfully applied to current-mode [22] and constant on-time (COT) [23] controlled Buck converters. It has also been extended to multiphase COT converters with phase overlapping [24, 25], and to passive-ripple architectures with exponentially decaying slopes [26]. To simplify the design process, equivalent circuit models based on Pade approximation have been proposed in [21, 22, 23]. Despite its impressive accuracy, the DF method provides limited physical insight, does not provide open-loop characterization, and involves cumbersome, topology-specific derivations.
Based on the sampled-data approach [27], a unified modeling method for various ripple-based control schemes is proposed in [28, 29]. By applying Shannon’s sampling theorem and infinite series summation, all sideband components are rigorously captured. Similar to the DF method, the sampled-data model remains accurate even beyond the switching frequency and does not rely on the small-ripple assumption. Additionally, it provides open-loop information and features a much simpler derivation process than DF-based models.
Although significant progress has been made in modeling analog-controlled converters, modeling efforts dedicated to digitally controlled converters remain comparatively limited. [29] extends the sampled-data approach to digitally controlled buck converters and derives an open-loop transfer function in the s-domain. Based on standard z-transform theory, [30] proposes a discrete-time model that accounts for the sampling and delay effects introduced by digital modulators, greatly simplifying digital compensator design.
Figure 1: PWM modulation schemes with possible synchronized ADC sampling instants (vertical solid, green) and modulating edges (vertical dashed, red) highlighted. (a) TEM, up-counting sawtooth carrier. (b) LEM, down-counting sawtooth carrier. (c) Symmetrical modulation, triangular carrier.
The influence of modulation strategies on system dynamics has also been explored. Fig.1 presents the modulation strategies. For example, [31] studies asymmetric carrier modulations and analyzes how leading-edge modulation (LEM) and trailing-edge modulation (TEM) affect loop stability and transient response in analog modulators, while [32] further investigates symmetric carrier modulation. [33] compares LEM and TEM in digital pulse-width modulators (DPWM) and proposes a second-order global equivalent circuit to facilitate controller design. [34] demonstrates that adopting LEM in a digitally controlled boost converter can eliminate the right-half-plane zero to enhance transient performance.
In addition, [35] and [36] show that different injection points and perturbation injection methods yield different loop gain measurements in digitally controlled converters. [36] further establishes the relationship between the responses obtained under different injection methods. Furthermore, [37] considers the case where the sampling rate is lower than the switching frequency and proposes a corresponding design methodology.
However, the impact of DPWM-ADC synchronization in digitally controlled converters has not been investigated in previous works. In practical implementations, the ADC sampling instant in digital buck converters is often dynamically aligned with the duty cycle, such that sampling occurs at the center of the on-time or off-time interval. This ensures the sampled inductor current approximates its average value under steady-state. However, such synchronization affects small-signal characteristics. The duty cycle perturbation affects the ADC samples. Despite its widespread use in engineering practice, to the authors’ best knowledge, this effect has not yet been accounted for in any existing small-signal model.
The remainder of this paper is organized as follows. Section II presents the small-signal sampled-data modeling under asymmetrical carrier modulation with explicit consideration of DPWM-ADC synchronization. A closed-form expression of the digital loop gain is derived using the modified z-transform, and it is shown that the analog loop gain can be obtained from the digital loop gain, thereby avoiding complicated infinite-series evaluations. Although the derivation begins with trailing-edge modulation (TEM), the conclusions are also extended to leading-edge modulation (LEM). Section III develops the small-signal model for the case of a symmetrical carrier. Section IV compares the proposed model with the simulation results and existing models, demonstrating its superior accuracy. Section V provides experimental validation, confirming that the proposed model exhibits good agreements in all cases. Section VI concludes the paper.
Throughout this paper, lowercase variables with a time argument (e.g. $i_{L}(t)$), denote time-domain signals that include both the periodic steady-state component (denoted by uppercase variables with a time argument, e.g. $I_{L}(t)$) and the small-signal perturbation (denoted by a hat, e.g. $\hat{i}_{L}(t)$). The DC value of a signal is indicated by an overline, e.g. $\overline{I_{L}}$.
II Small-Signal Modeling of Digital Buck Converters with Asymmetrical Carriers
Figure 2: A digital buck with ADC sampling at the center of the PWM On-interval and synchronized to the TEM DPWM.
Fig. 2 illustrates a digitally controlled synchronous Buck converter employing current-mode control and operating in forced continuous conduction mode (FCCM). The digital counter generates an up-counting sawtooth carrier $CNTR(t)$ with an amplitude of $CNTR_{MAX}$. When $CNTR$ reaches zero, the counter updates the shadow register. The shadow register serves to suppress spurious PWM pulses during abrupt changes in the duty-cycle command. The DPWM output will be high if the shadow register output is higher than $CNTR$ and will be low elsewhere. This configuration implements a TEM DPWM, in which the PWM rising edge is fixed at the beginning of each cycle, while the falling edge varies and moves according to the duty-cycle command.
To simplify the analysis, we define the beginning of each cycle as the steady-state ADC sampling instant ($kT_{S}$). The block diagram and timing are presented in Fig.2 and Fig.3, where $H_{i}$ is the current sensor gain and $T_{s}$ is the switching period. In practice, $H_{i}$ also includes the ADC gain. The scaled inductor current is denoted as $i_{s}$. The duty-cycle to the inductor current transfer function and the digital PI compensator transfer function[29] are:
$G_{id}(s)=V_{IN}\bigg[sL_{f}+R_{L}+(\frac{1}{sC_{f}}+R_{C})\parallel R_{LD}\bigg]^{-1}$ (1)
$G_{C}(z)=K_{P}+\frac{K_{i}T_{s}}{1-z^{-1}}$ (2)
Figure 3: Discrete-domain waveforms of a TEM buck converter. Dense and sparse vertical dashed lines denote the sampling instants at steady state and under perturbation, respectively. The ADC samples at integer multiples of $T_{S}$.
The ADC and shadow register are modeled as two sets of sample-and-hold blocks. Under the small-signal assumption, the DPWM comparator behaves as another ideal sampler, allowing the duty-cycle perturbation $\hat{d}$ to be treated as an impulse train [28]. When a sampling-hold-sampling sequence operates at the same sampling rate, it can be equivalently represented by a single sampler followed by a delay equal to the time difference between the two sampling instants. By applying this lemma twice, the 3 ideal samplers and 2 ZOH blocks found above can be simplified to a single sampler followed by a delay $T_{D}$, without any ZOH. Under the small-signal assumption, $T_{D}$ is defined as the steady-state time between the ADC sampling instant and the moment when the new calculated duty-cycle command takes effect.
Since DPWM-ADC synchronization guaranties that sampling occurs exactly at the midpoint of the PWM on-interval, as shown in Fig.4, in each cycle, the ADC sampling instant $t_{smp}$ is related to the on-time perturbation $\hat{t}_{on}$ of the previous cycle:
Figure 4: Sampling instant variation due to DPWM-ADC synchronization. Dense and sparse vertical dashed lines represent sampling points under steady-state and perturbed, respectively.
$t_{smp\_k}=kT_{S}+0.5\hat{t}_{on}[k-1]$ (3)
Under the small-signal assumption, the k-th sampled inductor current $i_{L}(t_{smp\_k})$ can be approximated its 1st-order Taylor expansion as:
$i_{s}(t_{smp\_k})\approx i_{s}(kT_{S})+\frac{di_{s}(t)}{dt}\bigg|_{t=kT_{S}}\times\frac{\hat{t}_{on}[k-1]}{2}$ (4)
where $i_{s}(kT_{S})$ is the scaled inductor current sampled at fixed instants $kT_{S}$, $I_{s}(kT_{S})$ is the steady-state sampled inductor current, which equals the average inductor current $\overline{I_{L}}$:
$i_{s}(kT_{S})=i_{s}[k]=I_{s}(kT_{S})+\hat{i}_{s}[k]=\overline{I_{L}}H_{i}+\hat{i}_{s}[k]$ (5)
Extracting perturbation terms from (4) and (5) yields:
$\hat{i}_{smp}[k]=\hat{i}_{s}[k]+\frac{di_{s}(t)}{dt}\bigg|_{t=kT_{S}}\times\frac{\hat{t}_{on}[k-1]}{2}$ (6)
Therefore, with (6) and Lemma 1, Fig.2 can be simplified to Fig.5. $G_{CM}(s)$ represents the analog feedback-to-modulator output transfer function and accounts for the side-band effects due to sampling, the ”CM” subscript stands for controller and modulator. $G_{Plant}(z)$ denotes the plant seen by the digital PI compensator $G_{C}(z)$. The transfer function of the pure discrete part in Fig. 5, from $\hat{i}_{s}[k]$ to $\hat{t}_{on}[k]$, can be derived as:
$G_{D}(z)=\frac{-G_{C}(z)\frac{T_{S}}{CNTR_{MAX}}}{1-[-G_{C}(z)\frac{T_{S}}{CNTR_{MAX}}]\times H_{sync}(z)}$ (7)
where the feedthrough due to DPWM-ADC alignment is:
$H_{sync}(z)=\frac{z^{-1}}{2}\frac{di_{L}(t)}{dt}\bigg|_{t=kT_{S}}\times H_{i}$ (8)
According to the sampling theorem, the impulse train $\hat{i}_{s}[k]$’s s-domain representation $\hat{i}_{s}^{*}(s)$ is:
$\hat{i}_{s}^{*}(s)=\hat{i}_{s}^{*}(s+jn\omega_{S})=\frac{1}{T_{s}}\sum_{n=-\infty}^{\infty}\hat{i}_{s}\!\left(s+jn\omega_{s}\right)$ (9)
isolating the $n=0$ case of the infinite summation gives:
$\hat{i}_{s}^{*}(s)=\frac{1}{T_{s}}\hat{i}_{s}(s)+\frac{1}{T_{s}}\sum_{\begin{subarray}{c}n=-\infty\\ n\neq 0\end{subarray}}^{\infty}\hat{i}_{s}\!\left(s+jn\omega_{s}\right)$ (10)
The second term of (10) models sideband couplings. From Fig.5, $\hat{i}_{s}(s)$ can be expressed as:
$\hat{i}_{s}(s)=\hat{t}_{on}^{*}(s)e^{-sT_{D}}G_{id}(s)H_{i}$ (11)
Since $\hat{t}_{on}[k]$ is also an impulse train, its s-domain representations are:
$\hat{t}_{on}^{*}(s)=\hat{t}_{on}^{*}(s+jn\omega_{S})=G_{D}(e^{sT_{S}})\hat{i}_{s}^{*}(s)$ (12)
Expanding the infinite summation term in (10) with (11) and then replacing $\hat{t}_{on}^{*}(s+jn\omega_{S})$ with $\hat{t}_{on}^{*}(s)$ according to (12) gives:
$\hat{i}_{s}^{*}(s)=\frac{1}{T_{s}}$ $\bigg[\hat{i}_{s}(s)+\hat{t}_{on}^{*}(s)\times$ (13)
$\sum_{\begin{subarray}{c}n=-\infty\\ n\neq 0\end{subarray}}^{\infty}e^{-(s+jn\omega_{S})T_{D}}G_{id}(s+jn\omega_{S})H_{i}\bigg]$
Substituting (13) into (12) to eliminate $\hat{i}_{s}^{*}(s)$, then isolating $\hat{t}_{on}^{*}(s)$ and $i_{s}(s)$ gives:
$\frac{\hat{t}_{on}^{*}(s)}{\hat{i}_{s}(s)}=\frac{\frac{G_{D}(z)}{T_{S}}}{1-\frac{G_{D}(z)}{T_{S}}\displaystyle\sum_{\begin{subarray}{c}n=-\infty\\ n\neq 0\end{subarray}}^{\infty}e^{-(s+jn\omega_{S})T_{D}}G_{id}(s+jn\omega_{S})H_{i}}$ (14)
$\quad\quad\quad\text{where: }\quad z=\exp(sT_{S})$
Hence, $G_{CM}(s)$ can be derived from (14) as:
$G_{CM}(s)\triangleq-\frac{\hat{d}(s)}{\hat{i}_{s}(s)}=-e^{-sT_{D}}\frac{\hat{t}_{on}^{*}(s)}{\hat{i}_{s}(s)}$ (15)
Figure 5: Block diagram of TEM digital buck without ZOH.
The analog loop gain $T_{i}$ can be expressed as:
$T_{i}(s)=G_{CM}(s)G_{id}(s)H_{i}$ (16)
From Fig. 5, the plant transfer function seen by the digital compensator can be easily found as:
$G_{Plant}(z)=\frac{T_{S}}{CNTR_{MAX}}\bigg[H_{sync}(z)+G_{MZ}(z)\bigg]$ (17)
where:
$G_{MZ}(z)=\frac{1}{T_{S}}\displaystyle\sum_{\begin{subarray}{c}n=-\infty\end{subarray}}^{\infty}\big[e^{-sT_{D}}G_{id}(s)H_{i}\big]_{s->s+jn\omega_{S}}$ (18)
The digital loop gain $T_{pul}$ is then:
$T_{pul}(z)=G_{Plant}(z)G_{C}(z)$ (19)
The loop gain measurement reads differently as the injection point changes [35, 36]. The relationship between $T_{i}$ and $T_{pul}$ is found to be:
$T_{i}(s)=\frac{T_{0}(s)}{1+T_{pul}(\exp(sT_{S}))-T_{0}(s)}$ (20)
where:
$T_{0}(s)=G_{c}(e^{sT_{S}})\frac{1}{CNTR_{MAX}}e^{-sT_{D}}G_{id}(s)H_{i}$ (21)
(20) agrees with the conclusion of [36] and provides a convenient method to evaluate the analog loop gain from the digital loop gain. Furthermore, (18) resembles the modified Z-transform[38] of $G_{id}H_{i}$ with a sampling delay of $T_{D}$. One can perform partial fraction on $G_{id}H_{i}$ and then utilize the following modified Z-transform identity to evaluate (18):
$\mathcal{Z}_{m}\!\left\{\sum_{r=1}^{\mathrm{PoleCount}}\frac{n_{r}}{s+d_{r}},\,T_{p}\right\}=\sum_{r=1}^{\mathrm{PoleCount}}\frac{n_{r}\,e^{d_{r}T_{p}}}{z\,e^{d_{r}T_{S}}-1}.$ (22)
where the delay $T_{p}$ must satisfy $T_{p}\in(0,\,T_{S})$, i.e., strictly within one sampling period. If the desired sampling delay $T_{D}$ is not in this range, it can be decomposed as:
$T_{D}=kT_{S}+T_{p},\quad k\in\mathbb{Z},\;T_{p}\in(0,\,T_{S}).$ (23)
In this case, the modified $z$-transform of a delay $T_{D}$ can be obtained by multiplying $z^{-k}$ with the modified $z$-transform of a delay $T_{p}$:
$G_{MZ}(z)=\mathcal{Z}_{m}\{G_{id}H_{i},\,T_{D}\}=z^{-k}\mathcal{Z}_{m}\{G_{id},\,T_{p}\}H_{i}.$ (24)
where in the case of TEM, $k=1$ and $T_{p}=0.5DT_{S}$, as shown in Fig.2 and Fig.3.
The above analysis for TEM can be extended to LEM. Table I summarizes the effective sampling delay $T_{D}$ under different modulation modes and sampling strategies.
TABLE I: $T_{D}$ Under Various Asymmetrical Modulations and Different Sampling Positions
Modulation + Sampling Position k $T_{p}$
TEM, Sample at the on-interval center 11 $0.5DT_{S}$
TEM, Sample at the off-interval center 0 $0.5(1+D)T_{S}$
LEM, Sample at the on-interval center 0 $(1-0.5D)T_{S}$
LEM, Sample at the off-interval center 11 $0.5(1-D)T_{S}$
By applying (24) to (17), the plant transfer function seen by the digital compensator can be obtained for both TEM and LEM cases, which provides the basis for loop compensation design, e.g., using the pole-placement method.
Figure 6: Key discrete-domain waveforms of a buck converter with a symmetrical carrier. Dense and sparse vertical dashed lines denote the sampling instants at steady state and under perturbation, respectively. The ADC samples at integer multiples of $T_{S}$.
III Small-Signal Modeling of Digital Buck Converters with Symmetrical Carriers
The above analysis for asymmetrical carriers can be extended to the case where the carrier is a symmetric triangular waveform. Fig.6 illustrates the key waveforms of a digital buck converter with a triangular carrier. The ADC samples when $CNTR$ reaches 0, which corresponds to the center of the PWM on-interval. Afterwards, the new duty cycle command is loaded into the shadow register when $CNTR$ reaches $CNTR_{MAX}$. Compared to the asymmetrical case, two important differences arise. First, in the symmetrical case, the ADC sampling instant is always stationary with respect to the carrier and does not vary with any perturbation. Consequently, there will be no $H_{\text{sync}}(z)$. Second, each impulse $\hat{t}_{on}[k]$ now maps to two impulses in $\hat{d}$ that have different delays ($T_{D1}$ and $T_{D2}$), which is similar to the sampled-data COT model in [29].
In this case, the plant transfer function in the $z$-domain seen by the digital compensator is given by:
$G_{\text{Plant,SYM}}(z)=\frac{T_{s}}{2\,\text{CNTR}_{\max}}\bigg($ $\mathcal{Z}_{m}\{G_{id},T_{D1}\}+$ (25)
$\mathcal{Z}_{m}\{G_{id},T_{D2}\}\bigg)H_{i}$
where $T_{D1}$ and $T_{D2}$ are given in Table II. This table also accounts for the alternative configuration where sampling occurs at the off-interval center, i.e., when $CNTR=\text{CNTR}_{\max}$ and the shadow register is updated at $CNTR=0$.
TABLE II: $T_{D1}$ and $T_{D2}$ Under Symmetrical Modulation and Different Sampling Positions
Sampling Position $T_{D1}$ $T_{D2}-T_{D1}$
Sample at the on-interval center $(1-0.5D)T_{S}$ $DT_{S}$
Sample at the off-interval center $0.5(1+D)T_{S}$ $(1-D)T_{S}$
Equation (25) provides the plant transfer function for digital compensator design. The open-loop transfer function in the digital domain can then be computed from (19), while the corresponding analog-domain loop gain can still be obtained directly using (20). In this case, $T_{0}(s)$ is modified as:
$T_{0,\text{SYM}}(s)=\frac{G_{c}(e^{sT_{s}})}{\text{2CNTR}_{\max}}(e^{-sT_{D1}}+e^{-sT_{D2}})G_{id}(s)H_{i}$ (26)
IV Comparison Between Existing Small-Signal Models and the Proposed Analytical Model
For a fair comparison, all models are evaluated under the same set of circuit and control parameters, summarized in Table III.
TABLE III: Parameters Used for Model Comparison
Symbol Description Value
$V_{\text{IN}}$ Input voltage $12V$
$D$ Steady-state duty cycle $27.596\%$
$T_{S}$ Switching period $10\mu s$
[…]
$R_{L}$ DC winding resistance (DCR) of $L_{f}$ $1m\Omega$
[…]
$R_{C}$ Equivalent series resistance (ESR) of $C_{f}$ $10m\Omega$
$R_{LD}$ Load resistance $0.33\Omega$
$CNTR_{MAX}$ Digital counter upper bound $1.2$
KpKp Compensator proportional gain $0.2$
KiKi Compensator integration gain 3142031420
[…]
Fig.9 presents the plant transfer function as perceived by the digital compensator, predicted by the proposed model, Dragan’s purely discrete-time model [30], and SIMPLIS simulations. It is observed that Dragan’s model, which neglects DPWM-ADC synchronization, consistently deviates from the simulation results. In contrast, the proposed model achieves an excellent match across the entire frequency range, validating its superior accuracy.
Fig. 10 illustrates the analog and digital loop gains using different perturbation methods. As reported in [35, 36], the two measurements exhibit close agreement in the mid-frequency range (approximately 100 Hz to half the switching frequency), but significant discrepancies appear at both low and high frequencies. At low frequencies, the digital loop gain behaves as an ideal integrator, theoretically yielding an infinite DC gain, whereas the analog loop gain settles to a finite value. This property can be rigorously verified by evaluating the limits of (20) and (19) as $s\to 0$.
To further verify the effectiveness of the proposed model, Fig. 11 compares its predictions against SIMPLIS simulations under symmetrical modulation, where the ADC sampling is performed at the turn-off center.
Table IV compares the proposed model with several established modeling approaches, including the sampled-data model in [29], the discrete-time model in [30], and the HTF-based model in [35]. Unlike the existing methods, the proposed model simultaneously incorporates DPWM-ADC synchronization, supports all three DPWM carrier types, and remains computationally inexpensive with low derivation complexity.
TABLE IV: Comparison between the proposed model and existing models of digital DC-DC converters
Model
Consider
DPWM-ADC
Alignment
Perspective Modulation Complexity
Yan’s[29] No Analog TEM Moderate
Dragan’s[30] No Digital TEM Low
Lin’s[35] No Analog TEM High
This work Yes
Analog &
Digital
TEM, LEM,
Symmetrical
Low
Figure 7: Analog loop gain (TEM + turn-off-centered sampling): the proposed model, Yan’s model [29], and simulation.
Figure 8: Analog closed-loop transfer function (TEM + turn-off-centered sampling): the proposed model, Yan’s model [29], and simulation.
Figure 9: Plant transfer function seen by the digital compensator (TEM + turn-off-centered sampling): the proposed model, Dragan’s model [30], and simulation.
Figure 10: Analog and digital loop gains predicted by the proposed model (TEM + turn-off-centered sampling).
Figure 11: Analog and digital loop gains under symmetrical triangular modulation with turn-off centered sampling: comparison of the proposed model and simulation.
V Experimental Verification
To further validate the proposed model, a synchronous buck converter prototype was built as shown in Fig. 12. The controller is implemented on a floating-point DSP (TMS320F28379) running at 200 MHz. Both the switching frequency and ADC sampling rate are set to 100 kHz. A high-resolution PWM is employed such that the PWM resolution far exceeds the 12-bit ADC resolution, thus avoiding limit cycles caused by duty-cycle quantization. Because the resolutions of DPWM and ADC are sufficiently high, quantization effects are negligible in the digital controller operation.
Figure 12: Experiment setup for loop gain measurements.
Figure 13: Plant characterization setup. $v_{p}(t)$ is the network analyzer AC output.
To obtain an accurate plant model for validation, the plant transfer function $G_{id}(s)H_{i}$ is experimentally identified using a network analyzer (Bode 100), as illustrated in Fig. 13. The Bode 100 generates an AC perturbation $v_{p}(t)$, which is combined with a DC bias through a resistive network and sampled by the ADC. After offset and scaling in the DSP, this perturbation is superimposed on the steady-state duty-cycle command, producing a perturbed PWM signal $d(t)$. The Bode 100 measures the transfer function from $d(t)$ to $i_{s}(t)$. Ideally, this transfer function is
$G_{id}(s)H_{i}=\frac{i_{s}(s)}{d(s)}=\frac{V_{IN}H_{i}}{V_{D}\times Z_{RLC}(s)},$ (27)
where $V_{D}$ is the amplitude of $d(t)$ and $Z_{RLC}(s)$ is the RLC impedance seen from the switching node $v_{SW}$. Since the open-loop identification is performed on the actual hardware ($V_{IN}=12$, $D=0.256$), all relevant parasitics (including MOSFET on-resistances and current-dependent inductor core losses) are inherently captured. Therefore, the identified transfer function in (28) is used in all subsequent comparisons, so that any discrepancy primarily reflects the accuracy of the proposed small-signal model rather than uncertainties in component values or operating conditions.
$G_{id}(s)H_{i}=12\times\frac{243147.4+429959502}{s^{2}+16837.45s+601451266}\times 0.1$ (28)
In the closed-loop experiments, the converter setup operates under TEM modulation with turn-on centered sampling. PI controller gains are set as $K_{p}=0.2$ and $K_{i}=0.2$. Fig. 14 shows the measured analog loop gain. The analog loop gain is obtained via Bode 100, using an injection transformer to insert perturbations as a voltage source between the ADC input and the low-side current-sensing resistor. Two input channels of the Bode 100 measure the voltages at both ends of the injection transformer relative to the ground.
Fig. 15 presents the comparison of the digital loop gain predicted by the proposed model against measurements obtained via two distinct injection methods. In the first method, akin to that in [35], the DSP samples the perturbation signal from the Bode 100 analyzer, combines it in software with the sampled inductor current, and then outputs both the pre- and post-injection current samples via DACs back to the Bode 100. The second method implements a Software-based Frequency Response Analyzer (SFRA) directly on the DSP: the SFRA injects discrete sinusoidal perturbations in software at selected frequencies and applies a Discrete-Time Fourier Transform (DTFT) to extract the magnitude and phase response at each frequency. This procedure is repeated sequentially over the frequency range to construct the bode plot. Fig. 15 demonstrates that both measurement methods yield equivalent results and closely agree with the model predictions.
The SFRA method enables digital-loop measurement without the need for additional hardware, and when combined with the highly accurate discrete-time model proposed in this work, it provides an effective framework for digital compensator design under real operating conditions.
Figure 14: Bode diagram of the calculated and measured analog loop gain.
Figure 15: Bode diagram of the calculated and measured digital loop gain.
VI Conclusions
This paper has presented an exact small-signal model for digitally controlled buck converters operating in FCCM under constant-frequency current-mode control, with explicit consideration of DPWM-ADC synchronization. Based on sampled-data theory, the proposed model rigorously captures all sideband components and provides accurate predictions of both analog and digital loop gains over a wide frequency range, including frequencies beyond the switching and sampling frequencies. This work further shows that, by using the modified $z$-transform, the digital loop gain can be computed in an accurate yet low-complexity closed form, and the analog loop gain can then be directly derived from the digital loop gain, thereby avoiding computationally intensive infinite-series evaluations.
The modeling procedure first considered asymmetrical sawtooth carriers (TEM and LEM) and was then extended to symmetrical triangular carriers. For each case, the sampling-to-PWM actuation delays were summarized in compact expressions that are convenient for control design. For asymmetrical carriers, the proposed analysis, for the first time, identifies and explains an additional feedthrough path introduced by DPWM-ADC synchronization, arising from the fact that the sampling instant moves with duty-cycle perturbations. In contrast, under symmetrical carriers the sampling instant remains fixed with respect to the carrier, no such feedthrough path exists, and the resulting loop dynamics are fundamentally different.
Compared with existing models, the proposed framework, for the first time, systematically treats all practically relevant combinations of carrier types (asymmetrical and symmetrical) and sampling instants in digitally controlled buck converters. In all cases examined, the analytical predictions of loop gains agree well with SIMPLIS simulations in both the analog and digital domains. The experimental results further confirm the accuracy and practical relevance of the proposed model. Moreover, the software-based SFRA yields digital loop-gain measurements that closely match both the proposed model and the network-analyzer-based results, indicating that SFRA can be effectively used to support digital controller design even in the absence of a dedicated network analyzer.
The combination of an analytically tractable model and experimentally validated accuracy makes the proposed approach a useful tool for the design and stability assessment of digitally controlled power converters with synchronized sampling. Future work will focus on extending the methodology to other converter topologies and multi-phase architectures.